OpenStudy (anonymous):

There are 10 members of a committee. One will be chosen as President, one as Vise President, one as Treasurer, and one as Secretary. In how many different ways can those offices be filled if the same person cannot serve in more than one office at the same time? a) 10!/4!6! b) 10!/4! c) 10! d) 10!/6!

OpenStudy (anonymous):

10! / 4 !

OpenStudy (anonymous):

How do you figure?

OpenStudy (anonymous):

there are 10 ways to choose president. for each choice of president, there are 9 choices left for vice president, then 8 choices left for vice president ... 10*9*8*7

OpenStudy (anonymous):

oh it should be d)

OpenStudy (anonymous):

nPr = n! / ( n-r)!

OpenStudy (anonymous):

10*9*8*7 = 10*9*8*7 * 6! / 6!

OpenStudy (anonymous):

i think maybe no.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oops you are right sorry

OpenStudy (anonymous):

10*9*8*7

OpenStudy (anonymous):

it is just that \[\frac{10!}{6!}\] is a silly way to write that

OpenStudy (anonymous):

my mistake laplace is right and i am wrong